# Mitigating Long-Tail Bias in HOI Detection via Adaptive Diversity Cache

### v1

@@ 4.3 Comparison with State-of-the-Art
4.4 Ablation Studies
4.4.1 Ablation Study of ADC Components
@@
To systematically validate the effectiveness of the proposed components and analyze their individual contributions, we conduct ablation experiments on each module. As shown in Tab. 2, starting from a baseline configuration (ADA-CM), we progressively add components and evaluate results across all metrics. From the experimental results, it can be observed that each additional component brings consistent improvements, with the complete method achieving optimal performance across all metrics. This effectiveness stems from addressing the fundamental issue of insufficient representative samples for rare interactions during training: CJCS expands the reference pool beyond the limited training set by caching high-quality diverse examples; in FACA concentrates cache capacity where maximum impact is achieved while reasonably constraining the cache capacity range, and provides additional gains through enhanced knowledge transfer. Each component complements the others, with their combination yielding the highest performance across all metrics.
4.4.2 Discussion of Cache Capacity in ACA
@@
4.4.3 Hyperparameter Sensitivity Analysis
@@
To understand the sensitivity of hyperparameters in final logits computation, we conduct ablation studies on the scaling factor and temperature parameter , as presented in Fig. 7. The experimental results reveal that the scaling factor demonstrates a clear trade-off between rare and non-rare category performance, with the highest full mAP achieved at , and the temperature parameter controls the sharpness of attention distribution over cache entries, with optimal performance achieved at .
@@
These observations highlight critical considerations for logits computation. The trade-off observed with indicates that excessive cache influence can overwhelm the model’s ability to process well-represented categories effectively, while insufficient cache contribution fails to adequately support rare category detection. The temperature parameter requires careful tuning to balance attention distribution: overly uniform attention fails to prioritize relevant cache entries, while excessively sharp focus may lead to over-reliance on limited historical information. The optimal configuration should achieve the best balance between leveraging historical context and maintaining prediction flexibility across different category types.
4.5 Zero-shot Performance Analysis
@@
To evaluate ADC’s effectiveness in zero-shot scenarios, we conduct experiments with different baseline methods on HICO-DET. As shown in Tab. 3, in terms of RF-UC and NF-UC settings, our approach marks a significant advancement, particularly for the performance of unseen categories. Relative to the state-of-the-art method EZ-HOI [lei2024ez], our unseen accuracy achieves a consistent gain of 0.43% and 0.80% for RF-UC and NF-UC, respectively. This is because the baseline model’s zero-shot design enables the cache to accumulate reliable interaction patterns even for rare/unseen categories, which ADC can effectively leverage. Moreover, ADC universally benefits all baseline methods, achieving consistent improvements across the board. Notably, beyond the gains in unseen categories, ADC also simultaneously enhances both Full and Seen mAP, demonstrating its comprehensive effectiveness.
@@
| Method | Setting | Full | Seen | Unseen |
| GEN-VLKT [liao2022gen] | RF-UC | 30.56 | 32.91 | 21.36 |
| HOICLIP [ning2023hoiclip] | RF-UC | 32.99 | 34.85 | 25.53 |
| CLIP4HOI [mao2023clip4hoi] | RF-UC | 34.08 | 35.48 | 28.47 |
| LOGICHOI [li2023neural] | RF-UC | 33.17 | 34.93 | 25.97 |
| ADA-CM [lei2023efficient] | RF-UC | 33.01 | 34.35 | 27.63 |
| + ADC | RF-UC | 38.17 | 38.73 | 35.96 |
| - | +5.16 | +4.38 | +8.33 | |
| HOIGen [guo2024unseen] | RF-UC | 33.86 | 34.57 | 31.01 |
| + ADC | RF-UC | 35.56 | 35.78 | 34.67 |
| - | +1.70 | +1.21 | +3.66 | |
| EZ-HOI [lei2024ez] | RF-UC | 36.73 | 37.35 | 34.24 |
| + ADC | RF-UC | 37.16 | 37.39 | 37.06 |
| - | +0.43 | +0.04 | +2.82 | |
| GEN-VLKT [liao2022gen] | NF-UC | 23.71 | 23.38 | 25.05 |
| HOICLIP [ning2023hoiclip] | NF-UC | 27.75 | 28.10 | 26.39 |
| CLIP4HOI [mao2023clip4hoi] | NF-UC | 28.90 | 28.26 | 31.44 |
| LOGICHOI [li2023neural] | NF-UC | 27.95 | 27.86 | 26.84 |
| ADA-CM [lei2023efficient] | NF-UC | 31.39 | 31.13 | 32.41 |
| + ADC | NF-UC | 36.73 | 37.17 | 34.95 |
| - | +5.34 | +6.04 | +2.54 | |
| HOIGen [guo2024unseen] | NF-UC | 33.08 | 32.86 | 33.98 |
| + ADC | NF-UC | 36.01 | 36.47 | 34.19 |
| - | +2.93 | +3.61 | +0.21 | |
| EZ-HOI [lei2024ez] | NF-UC | 34.84 | 34.47 | 36.33 |
| + ADC | NF-UC | 35.64 | 34.93 | 37.48 |
| - | +0.80 | +0.46 | +1.15 |
4.6 Qualitative Results

### v2

@@ IV-C Comparison with State-of-the-Art
IV-D Ablation Studies
IV-D1 Ablation Study of ADC Components
@@
To systematically validate the effectiveness of the proposed components and analyze their individual contributions, we conduct ablation experiments on each module. As shown in Tab. II, starting from a baseline configuration (ADA-CM), we progressively add components and evaluate results across all metrics. From the experimental results, it can be observed that each additional component brings consistent improvements, with the complete method achieving optimal performance across all metrics. This effectiveness stems from addressing the fundamental issue of insufficient representative samples for rare interactions during training: CJCS expands the reference pool beyond the limited training set by caching high-quality diverse examples; in FACA concentrates cache capacity where maximum impact is achieved while reasonably constraining the cache capacity range, and provides additional gains through enhanced knowledge transfer. Each component complements the others, with their combination yielding the highest performance across all metrics.
IV-D2 Discussion of Cache Capacity in ACA
@@
Furthermore, this stability in cache quality (Fig. 6) demonstrates the model’s inherent robustness to noisy pseudo-labels. As the average joint score progressively increases and stabilizes, the selection mechanism effectively filters out low-confidence entries. Simultaneously, as shown in the t-SNE visualization (Fig. 4), incorrect pseudo-labels typically appear as feature outliers in the latent space. These sparse outliers are statistically suppressed by the dominant signal from the dense, correct clusters during the affinity-based retrieval, ensuring that the final predictions remain robust even in the presence of initial noise.
IV-D3 Hyperparameter Sensitivity Analysis
@@
To understand the sensitivity of hyperparameters in the final logits computation, we conduct ablation studies on the scaling factor and temperature parameter , as presented in Fig. 7. The experimental results reveal that the scaling factor demonstrates a clear trade-off between rare and non-rare category performance, with the highest full mAP achieved at . The temperature parameter controls the sharpness of the attention distribution over cache entries, with optimal performance achieved at .
@@
These observations highlight critical considerations for logits computation. The trade-off observed with indicates that excessive cache influence can overwhelm the model’s ability to process well-represented categories effectively, while insufficient cache contribution fails to adequately support rare category detection. The temperature parameter can balance attention distribution: overly uniform attention fails to prioritize relevant cache entries, while excessively sharp focus may lead to over-reliance on limited historical information.
A critical concern for practical deployment is whether these hyperparameters require dataset-specific tuning. Our experiments confirm the robustness of ADC: All main results reported in Tab. I (across different backbones and datasets) use the identical hyperparameter setting (). This demonstrates that ADC does not require careful per-dataset tuning to achieve state-of-the-art performance. As shown in Fig. 7, although optimal values exist, ADC maintains consistent performance gains over the baseline across a wide range of parameter values. The performance does not collapse even with sub-optimal settings, confirming the method’s stability and ease of deployment.
| Method | RF-UC Setting | NF-UC Setting | ||||
| Full | Seen | Unseen | Full | Seen | Unseen | |
| GEN-VLKT [34] | 30.56 | 32.91 | 21.36 | 23.71 | 23.38 | 25.05 |
| HOICLIP [38] | 32.99 | 34.85 | 25.53 | 27.75 | 28.10 | 26.39 |
| CLIP4HOI [37] | 34.08 | 35.48 | 28.47 | 28.90 | 28.26 | 31.44 |
| LOGICHOI [30] | 33.17 | 34.93 | 25.97 | 27.95 | 27.86 | 26.84 |
| ADA-CM [27] | 33.01 | 34.35 | 27.63 | 31.39 | 31.13 | 32.41 |
| + ADC (Ours) | 38.17 | 38.73 | 35.96 | 36.73 | 37.17 | 34.95 |
| +5.16 | +4.38 | +8.33 | +5.34 | +6.04 | +2.54 | |
| HOIGen [15] | 33.86 | 34.57 | 31.01 | 33.08 | 32.86 | 33.98 |
| + ADC (Ours) | 35.56 | 35.78 | 34.67 | 36.01 | 36.47 | 34.19 |
| +1.70 | +1.21 | +3.66 | +2.93 | +3.61 | +0.21 | |
| EZ-HOI [26] | 36.73 | 37.35 | 34.24 | 34.84 | 34.47 | 36.33 |
| + ADC (Ours) | 37.16 | 37.39 | 37.06 | 35.64 | 34.93 | 37.48 |
| +0.43 | +0.04 | +2.82 | +0.80 | +0.46 | +1.15 | |
IV-E Zero-shot Performance Analysis
@@
To evaluate ADC’s effectiveness in zero-shot scenarios, we conduct experiments with different baseline methods on HICO-DET. As shown in Tab. III, in terms of RF-UC and NF-UC settings, our approach marks a significant advancement, particularly for the performance of unseen categories. Relative to the state-of-the-art method EZ-HOI [26], our unseen accuracy achieves a consistent gain of 0.43% and 0.80% for RF-UC and NF-UC, respectively. This is because the baseline model’s zero-shot design enables the cache to accumulate reliable interaction patterns even for rare/unseen categories, which ADC can effectively leverage. Moreover, ADC universally benefits all baseline methods, achieving consistent improvements across the board. Notably, beyond the gains in unseen categories, ADC also simultaneously enhances both Full and Seen mAP, demonstrating its comprehensive effectiveness.
@@
| Method | Backbone | SG1 | SG2 | SG3 |
| EZ-HOI [26] | R50 | 16.57 | 14.62 | 15.73 |
| + ADC | R50 | 17.64 | 15.53 | 16.42 |
| - | +1.07 | +0.91 | +0.69 | |
| ADA-CM [27] | R50 | 15.64 | 15.02 | 15.76 |
| + ADC | R50 | 17.16 | 16.93 | 17.01 |
| - | +1.52 | +1.91 | +1.25 | |
| HOIGEN [15] | R50 | 14.40 | 13.27 | 14.53 |
| + ADC | R50 | 16.27 | 15.66 | 16.28 |
| - | +1.87 | +2.39 | +1.75 |
IV-F Systematic Generalization Analysis
To further evaluate the model’s capability in handling the compositional nature of HOI detection—where objects and verbs may appear in novel combinations—we conduct experiments on the HICO-DET-SG (Systematic Generalization) splits [18]. Unlike standard random splits, SG splits are designed to test the model’s ability to recombine learned concepts for unseen triplets.
As shown in Tab. IV, ADC achieves consistent improvements across all three systematic generalization splits (SG1, SG2, SG3). Specifically, on the challenging SG1 split, ADC boosts the performance of EZ-HOI and ADA-CM by 1.07% and 1.52% mAP, respectively. This indicates that while our method focuses on mitigating long-tail bias via frequency-aware caching, the diversity mechanism in ADC effectively retrieves high-quality historical prototypes to “patch” unreliable predictions for rare or unseen compositions. This confirms that ADC not only addresses frequency imbalance but also enhances the model’s compositional generalization capability.
IV-G Efficiency and Online Adaptation Analysis
IV-G1 Computational Overhead
We analyze the computational cost of ADC in terms of inference time and memory usage. As presented in Tab. V, compared to the base models, ADC introduces a negligible memory overhead (e.g., increasing from 7619M to 8926M for EZ-HOI), as we only store lightweight feature vectors rather than heavy model parameters. Regarding time efficiency, while the retrieval mechanism introduces additional latency (approximately –), it remains far more efficient than standard gradient-based adaptation methods (which typically require 6–10 latency). This represents a superior trade-off between efficiency and the substantial accuracy gains observed.
IV-G2 Online Setting
We further evaluate ADC in a realistic Online setting, where the cache is built incrementally from scratch. As shown in Tab. V, we report results on both HICO-DET and V-COCO. Although the ”cold start” problem leads to slightly lower performance compared to the Offline setting, ADC still achieves robust improvements over the static baseline (+1.09 mAP), further verifying its effectiveness across different datasets. This confirms that our Confidence-Diversity Joint Cache Selection (CJCS) mechanism can effectively curate high-quality features on-the-fly.
| Method | HICO-DET (Online) | V-COCO (Online) | |||||
| mAP | Cost | mAP | Cost | ||||
| Full | Rare | Time | Mem. | Time | Mem. | ||
| EZ-HOI [26] | 38.33 | 36.70 | 1h 23m | 7619M | 60.50 | 23m | 6822M |
| + ADC | 38.81 | 39.57 | 2h 20m | 8926M | 60.97 | 39m | 8410M |
| ADA-CM [27] | 38.39 | 37.46 | 11m | 4886M | 58.50 | 8m | 4758M |
| + ADC | 39.48 | 40.60 | 39m | 6908M | 61.90 | 14m | 6916M |
| HOIGen [15] | 33.56 | 32.10 | 12m | 4612M | - | - | - |
| + ADC | 36.03 | 39.46 | 17m | 4920M | - | - | - |
IV-H Qualitative Results